OpenStudy (anonymous):

Need Pre-calc help Medal + fan for best helper Find the angle between the given vectors to the nearest tenth of a degree. u = <6, -1>, v = <7, -4>

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

We have some kind of handy formula for dealing with this, yes? \[\Large\rm \cos \theta=\frac{\vec u\cdot \vec v}{||u||~||v||}\]Where theta is the angle between them.

OpenStudy (anonymous):

and ||u|| is the magnitude?

zepdrix (zepdrix):

\[\Large\rm \cos \theta=\frac{<6, -1>\cdot<7, -4>}{||<6, -1>||~||<7, -4>||}\]Yes :) Length of the vector, magnitude.

OpenStudy (anonymous):

can you remind me how to find the magnitude? I'm very rusty at this stuff

zepdrix (zepdrix):